ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

12. Probability and Statistics

1991 MS Classification: 60F05 Geluk, Jaap, Erasmus University Rotterdam Second order regular variation and the domain of attraction The domain of attraction of stable distributions can be derived in a direct way using Fourier transforms and regular variation techniques. We characterize second order regular variation of the tail sum of a distribution function together with a balance condition on the tails in terms of the behavior of the charcteristic function near zero. Results are joint with L. De Haan and Liang Peng respectively.